Exercise 11-15 Dividend yield computation and interpretation LO A3 Company Annual Cash Dividend per Share Market Value per Share 1 $ 10.00 $ 163.93 2 7.00 72.92 3 8.50 80.19 4 1.80 118.15 Compute the dividend yield for each of these four separate companies.
Answer #1

Dividend yield = dividend per share/market value per share

1) Dividend Yield = 10/163.93 *100 = 6.1%

2) Dividend yield = 7/72.92 * 100 = 9.599 = 9.6%

3) Dividend yield = 8.5/80.19 *100 = 10.599 = 10.6%

4) Dividend yield = 1.8/118.15 * 100 = 1.52%